OpenStudy (anonymous):

This puzzles me There are two options: Option 1: Find a formula that is used in sports or leisure activities. Be sure that that formula has at least two variables. Rewrite your formula to solve for the other variable. Show all steps in your work. Option 2: Create a fictional formula of your own using at least four symbols (no numbers or letters). Try using the symbols above the number keys on your keyboard. Once you have created your formula, rewrite your formula to solve for one of the other symbols. Show all steps in your work.

OpenStudy (phi):

How about the number of points a team gets in football, assuming they score only td's or field goals ?

OpenStudy (anonymous):

Why make such an assumption? (;

OpenStudy (anonymous):

No clue math isnt my strong point and word problems have been my problems all day

OpenStudy (anonymous):

What's your favourite leisure activity?

OpenStudy (anonymous):

Playing videogames and eating ._.

OpenStudy (anonymous):

Any video games you play with a scoring system?

OpenStudy (anonymous):

Hmmm....League of Legends?

OpenStudy (anonymous):

I'm only vaguely familiar with LoL; is there any way score is calculated in-game from player kills, minion kills, etc? If so you could write an equation relating these.

OpenStudy (anonymous):

Yes actually but how would that be done?

OpenStudy (anonymous):

Well, let's say you get one point per minion kill and ten points per player kill. If that was the case, you could write: s = m + 10p where s is your score, m is your minion kill count, and p is your player kill count. You can, of course, adjust this with the actual values if there is a system like this, and add additional variables if you so desire.

OpenStudy (anonymous):

Hmmm nope not at all because they are not calculated by points like that but PFS games usually do like assists are worth 25 or 75 while the kill is 100 or so

OpenStudy (anonymous):

If you have an FPS game in mind with a scoring system like that you could use that instead.

OpenStudy (anonymous):

Alright so just set it up like you said?

OpenStudy (anonymous):

So its 75=a+3(25) with 25 being the points for the assists, 3 being the assists, and a being the kills while the 75 is the final score. Did I do it right?

OpenStudy (anonymous):

So I just need to make an equation to where I find the missing variable?

OpenStudy (anonymous):

Not quite. In the equation you've written, you're counting a kill as one point which I imagine you don't want. You also need to use variables for the score and number of assists, meaning you'll have an equation like the following: s = 100k + 25a where s = score, k = kills, a = assists

OpenStudy (anonymous):

You can then rearrange this equation to show, for example, the number of kills in terms of score and assists.

OpenStudy (anonymous):

Alright so I was kinda right I just forgot something

OpenStudy (anonymous):

Thanks <3